OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

y = 2x + 6 y = 1/2x + 3 y = 1/2x + 6 y = 2x + 3

OpenStudy (mathstudent55):

The slopes of perpendicular lines are negative reciprocals. That means if you know the slope of a line, the slope of its perpendiculars is the number that when you multiply by the slope you get -1. All this means that if you know the slope of a line, flip it and change the sign and you now have the slope of its perpendiculars.

OpenStudy (mathstudent55):

Since the given line is in slope-intercept form, y = mx + b, where m = slope, and b = y-intercept, if you compare y = 2x + 2, you can read the slope directly.